A certain weight of pure CaCO3 is made to react completely with 200 ml of a HCl solution to give 224 ml of CO2 gas at S.T.P. The normality of HCl solution is.....

Grade:12th pass

1 Answers

Pooja Aswani
6 years ago
According to the equation,CaCO3 + 2HCl -> CaCl2 + CO2 + H2O moles of HCl/2 =moles of CO2 moles of CO2=volume at stp(in L)/22.4=>0.224/22.4=0.01Moles of HCl=2×moles of CO2=>2×0.01=0.02Molarity =moles of solute/volume of solution(in L)=>0.02/0.2=0.1MNormality=molarity× n factor=>0.1×1=0.1N